Find the area of each circle. Round to the nearest tenth.
There is a drawing of a circle with a diameter of 21 ft. I divided the diameter in half to get a radius of 10.5. I then used the formula:A= pie r^.
A= pie r^
A= 3.14 (10.5)
A= 346.2ft
Did I do this right?

Found 3 solutions by unlockmath, Kyrashea21, richwmiller:
Answer by unlockmath(1688)

Hello,
Perfect job---Well done!:-)Just a couple points to watch--it's pi and the equation is area = pi radius squared
therefore you'll have square feet not feet.
RJ
Check out a new book I wrote at:
www.math-unlock.com

Answer by Kyrashea21(1)

Yes you did perfect.

Answer by richwmiller(17219)

Find the area of each circle. Round to the nearest tenth.
There is a drawing of a circle with a diameter of 21 ft. I divided the diameter in half to get a radius of 10.5. I then used the formula:A= pie r^.
A= pie r^
A= 3.14 (10.5)
A= 346.2ft
Did I do this right?
Almost.
pi*r^2
not apple pie
and 3.14*10.5^2 not 3.14*10.5
also 3.14*10.5^2=346.2 ft^2 not feet